OpenStudy (anonymous):

Cube 1 has sides of length L, has an area A, and a volume V. Cube 2 has an area that is 1/2 A. What is the length of the sides and the volume of Cube C2? How smart are you?

OpenStudy (anonymous):

to answer your second question first, i am dumb as a box of rocks

OpenStudy (anonymous):

to answer the second question, lets make the length of the first cube 1, so the area is 6 and the volume is 1 then the area of the second cube is 3, so the length of each side is \(\frac{1}{\sqrt{2}}\) and the volume is \(\frac{1}{4\sqrt{2}}=\frac{\sqrt{2}}{8}\)